OpenStudy (anonymous):

What is the sum of the solutions of ? 2|x -1|-4=-2 A. -2 B. 0 C. 1 1/2 D. 2

OpenStudy (anonymous):

2|x -1|-4=-2 /+4 2|x -1|=2 /:2 |x -1|=1 x-1=1 x=2 x-1=-1 x=0 2+0=2

OpenStudy (anonymous):

I actually had this one right then! Thank you so much!